OpenStudy (ffasinger):

What is the equation of the following line? Be sure to scroll down first to see all answer options.

OpenStudy (ffasinger):

A. y = -7x B. y = 21x C. y = 7x D. y = -3x E. y = 1/5x F. y = 3x

OpenStudy (anonymous):

So in this problem we know 2 points for sure, they are both 0 at the same time and when x=1, then y=7. So our y is growing at a rate of 7 * x. So y = 7x!

OpenStudy (ffasinger):

i have som more can you please help me with those too ill medal

OpenStudy (anonymous):

Of course I will do what I can, leaving in 30 minutes.

OpenStudy (ffasinger):

got it i will tag u